An impedance coil having a resistance of 15 ohms and an inductance of 0.08 henry is connected in series with a variable capacitor. If a 117 volt 60 cps source is connected to the circuit, what will be the capacitance of the capacitor and the current

(a) when the circuit has a power factor of 0.866 leading and (b) when the voltage drop across the coil is equal to the voltage drop across the capacitor?


1
Expert's answer
2021-03-08T03:08:25-0500

Explanations & Calculations


  • Consider the sketch attached




  • Knowns are

"\qquad\qquad
\begin{aligned}
\small R_L&= \small 15\Omega\qquad\qquad\qquad X_L=2\pi fL=30.16\Omega\\
\small \cos\theta&=\small 0.866\to\theta =30^0 (for\,part\,a\,only)
\end{aligned}"

  • All the resistance, inductance & capacitance are connected in series hence the same current flows through each.


a)

  • From the phasor diagram for part (a),

"\qquad\qquad
\begin{aligned}
\small V_R&= \small 117\cos\theta =117\times 0.866=101.32V
\end{aligned}"

  • Then by "\small V=iR",

"\qquad\qquad
\begin{aligned}
\small i&= \small \frac{101.32V}{15\Omega}=\bold{6.75A}
\end{aligned}"

  • Then the voltage drop across the inductive part is,

"\qquad\qquad
\begin{aligned}
\small V_L&= \small iX_L=203.58\,V
\end{aligned}"

  • Finally,

"\qquad\qquad
\begin{aligned}
\small 117\sin\theta&= \small V_c-V_L\\
\small V_c&= \small 117\sin30+203.58=262.08V\\
\small C&= \small \frac{6.75A}{2\pi\times 60Hz\times 262.08V}\\
&= \small \bold{68.32\,\mu F}
\end{aligned}"

b)

  • By refering to the phasor diagram for part (b),

"\qquad\qquad
\begin{aligned}
\small V_{coil}&= \small V_c\\
\small iZ_{coil}&= \small iX_c\\
\small \sqrt{R^2+X_L^2}&=\small X_c=\frac{1}{2\pi f C}\\
\small C&=\small \frac{1}{2\pi \times 60\times \sqrt{15^2+30.16^2}}\\
&= \small \bold{78.75\,\mu F}
\end{aligned}"

  • Then capacitive reactance is

"\qquad\qquad
\begin{aligned}
\small X_c&= \small\frac{1}{2\pi f C}\\
&= \small \frac{1}{2\pi\times 60\times 78.75\times10^{-6} }\\
&= \small 33.68 \,\Omega 
\end{aligned}"

  • FInally,

"\qquad\qquad
\begin{aligned}
\small 117^2 &= \small V_R^2+(V_c-V_L)^2\\
&=\small i^2\big[R^2+(X_c-X_L)^2\big]\\
\small i&= \small \frac{117}{\sqrt{15^2+(33.68-30.16)^2}}\\
&= \small \bold{7.59\,A}
\end{aligned}"